OpenStudy (anonymous):

sing complete sentences, compare the key features and graphs of sine and cosine. What are their similarities and differences?

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

Key features would be x-intercepts, y-intercepts, amplitude, and period

OpenStudy (anonymous):

thats it? is that the answer? :O

OpenStudy (misty1212):

i am not a very good singer ♩ ♪ ♫ ♬ they look pretty much the same ♩ ♪ ♫ ♬ ♩ ♪ ♫ ♬ except cosine of zero is one and sine of zero is zero ♩ ♪ ♫ ♬ ♩ ♪ ♫ ♬so cosine looks just like sine♩ ♪ ♫ ♬ ♩ ♪ ♫ ♬only shifted to the left \(\frac{\pi}{2}\) units ♩ ♪ ♫ ♬ ♩ ♪ ♫ ♬ which is another way of saying \(\sin(x+\frac{\pi}{2})=\cos(x)\)♩ ♪ ♫ ♬

OpenStudy (anonymous):

So like cos x y-int is at 0,1 and it has a x-intercept at pi/2 and 3pi/2 so the period would be 2pi for cos x and both cos x and sin x have the same amplitude and period

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

thank you!